OpenStudy (anonymous):

If the numerator of a fraction is increased by 3, the fraction becomes 3/4. If the denominator is decreased by 7, the fraction becomes 1. Determine the original fraction. Which of the following equations represents "If the numerator of a fraction is increased by 3, the fraction becomes 3/4"? (Hint: cross products) 3n + 9 = 4d 4n + 3 = 3d 4n + 12 = 3d

OpenStudy (anonymous):

start with \[\frac{n+3}{d}=\frac{3}{4}\] \[\frac{n}{d-7}=1\]